OpenStudy (clamin):

PLEASE HELP!! 3(2 - 8x) =-13-5x

OpenStudy (jessiegonzales):

3(2-8(1)=-12-5(1) x=1

OpenStudy (clamin):

i dont know where you got that 1from @jessiegonzales

OpenStudy (nikato):

remember distribute property a(b-c)= ab- ac so 3(2-8x)= 3(2) - 3(8x) = 6-24x so the equation becomes 6-24x=-13-5x +24x +24x Add 24x to both sides -------------- 6=-13+19x +13 +13 Add 13 to both sides ---------- 19=19x -- --- divied both sides by 19 19 19 x=1